Calculate the density of a proton, given that the mass of a proton is 1.0073 amu and the diameter of a proton is 1.72×10−15 m.
1
Expert's answer
2019-09-13T05:49:50-0400

Mass of proton = 1.0073amu = "1.0073\times1.66054\times10^{-27}\ kg=1.6726\times10^{-27 }\ kg"

Volume of proton = "\frac{4}{3}\pi({\frac{d}{2}})^3=\frac{4}{3}\pi(\frac{1.72\times 10^{-15}}{2})^3"

"=2.66\times 10^{-45}\ m^3"


Density "=\frac{mass}{volume}=\frac{1.6726\times 10^{-27}}{2.66\times10^{-45}}=6.288\times 10^{17}\ kg\ m^{-3}"